# Federated Prototype-based Contrastive Learning for Privacy-Preserving Cross-domain Recommendation

### v1

@@ III-B2 Clustering and Prototype Selection
Finally, we upload each domain’s representative prototype set $C^{i}$ and overlapping user set $O^{i}$ to the global server.
III-B3 Prototype-based Contrastive Learning
In the previous subsection, we uploaded representative prototypes to the global server. The server then aggregates these prototypes to derive both local and global prototypes (as further discussed in the next section) and subsequently returns them to local domains. Due to data privacy constraints, direct transfer of original data is not feasible. To address this problem, we employ prototypes and introduce prototype-based CL losses, which comprise a global part and a local part. These prototypes represent the collective preferences of multiple users rather than any single individual, making it more difficult for an attacker to associate a prototype with a specific user.
To transfer knowledge from other domains to domain $D^{i}$, we enforce the alignment of user embeddings with corresponding global prototypes while distancing them from distinct global prototypes. The global prototype-based CL loss is defined as follows:
$\small L_{global}^{i}=-log\frac{exp(f(\textbf{e}_{u}^{i},\textbf{g}_{k}^{i}))}{exp(f(\textbf{e}_{u}^{i},\textbf{g}_{k}^{i}))+\sum_{\textbf{g}_{j}^{i}\in A(\textbf{g}_{k}^{i}),j\not=k}exp(f(\textbf{e}_{u}^{i},\textbf{g}_{j}^{i}))},$ (6)
where $\textbf{g}_{k}^{i}$ denotes the global prototype corresponding to cluster $k$, to which the user embedding $\textbf{e}_{u}^{i}$ belongs. We regard $(\textbf{e}_{u}^{i},\textbf{g}_{k}^{i})$ that belongs to the same cluster as a positive pair. Conversely, $\textbf{g}_{j}^{i}$ denotes the global prototype corresponding to cluster $j$, to which user embedding $\textbf{e}_{u}^{i}$ doesn’t belong, and $(\textbf{e}_{u}^{i},\textbf{g}_{j}^{i})$ forms a negative pair. $A(\textbf{g}_{k}^{i})$ is the set of global prototypes excluding $\textbf{g}_{k}^{i}$. $f$ indicates a similar function. We define it as:
$f(\textbf{e}_{u}^{i},\textbf{g}_{k}^{i})=\frac{\textbf{e}_{u}^{i}\cdot\textbf{g}_{k}^{i}}{||\textbf{e}_{u}^{i}||\times||\textbf{g}_{k}^{i}||}/\tau,$ (7)
where $\tau$ represents the temperature coefficient, which controls the concentration strength of representation [36].
In addition to the global prototype-based CL loss, we introduce the local prototype-based CL loss to align $\textbf{e}_{u}^{i}$ with local prototypes of each domain through domain-wise CL in the latent space, enhancing inter-domain knowledge sharing. The local prototype-based CL loss is defined as follows:
@@ III-B3 Prototype-based Contrastive Learning
The global and local prototypes capture cluster-relevant information at different granularity, guiding the transfer of user interests from various perspectives.
III-B4 Local Training
After refining the user embedding $\textbf{e}_{u}^{i}$, we concatenate it with the item embedding $\textbf{e}_{v}^{i}$ and feed them into MLP layers to predict the user’s preferences. We aim to minimize the following loss function:
@@ III-B4 Local Training
where $l$ denotes the cross-entropy loss function, $\hat{r}^{i}$ is the predictive label, and $r^{i}$ is the ground-truth label.
The total loss function is defined as follows:
@@
where $\alpha$ is the trade-off parameter that balances the prediction loss and prototype-based CL losses. The detailed training process is in Algorithm 1 of Appendix A.
III-C Global Server Aggregation
After receiving representative prototype set $C$ and overlapping user set $O$ from all domains, the global server calculates global prototypes for each domain. First, for overlapping user set $o_{k}^{i}\in O^{i}$, we construct the prototype set $\hat{C}_{k}^{i}$ that includes all representative prototypes containing overlapping users with $o_{k}^{i}$ across domains:
$\hat{C}_{k}^{i}=\bigcup_{{i^{\prime}}\in D,{k^{\prime}}\leq{K^{\prime}}}\{\textbf{c}_{k^{\prime}}^{i^{\prime}}|o_{k^{\prime}}^{i^{\prime}}\cap o_{k}^{i}\not=\emptyset\}.$ (11)
@@ III-C Global Server Aggregation
$\textbf{g}_{k}^{i}=\frac{1}{\hat{K}}\sum_{k=1}^{\hat{K}}\hat{\textbf{c}}_{k}^{i},$ (12)
where $\hat{K}$ is the length of prototype set $\hat{C}^{i}_{k}$.
@@
Different from the global prototypes, for the local prototypes, we select some representative prototypes in all domains rather than a single one via similarity calculation. Specifically, for each representative prototype $\textbf{c}_{k}^{i}$, we first calculate the cosine similarity between $\textbf{c}_{k}^{i}$ and representative prototypes from other domains that include overlapping users with $o_{k}^{i}$. Then, we select the representative prototype with maximum similarity to $\textbf{c}_{k}^{i}$ in each domain to form the local prototype set $L_{k}^{i}$ for cluster $k$:
@@
III-D Privacy Preserving Analysis
@@
The proposed FedPCL-CDR effectively ensures user privacy. Firstly, within the FL framework, the data in each domain remains private and localized, significantly reducing the risk of user privacy leakage [37, 38]. Secondly, knowledge transfer across domains is facilitated through prototypes, which naturally protect data privacy [39]. These prototypes are 1-D vectors derived from averaging low-dimensional representations of samples within the same cluster, which is an irreversible process. We provide a detailed example in Appendix B.
IV Experiments
IV-A Experimental Settings
IV-A1 Datasets.
@@ IV-A Experimental Settings
Motivated by CDR methods [14, 19, 8], we conduct experiments on four real-world benchmark subsets from the Amazon dataset [^1] https://cseweb.ucsd.edu/ jmcauley/datasets/amazon/links.html: Phone, Electronics (Elec), Clothing, Shoes nad Jewelry (Cloth), and Sport, and three subsets from the Douban dataset [^2] https://www.dropbox.com/s/u2ejjezjk08lz1o/Douban.tar.gz?e=2&dl=0: Book, Movie, and Music. We combine them into four CDR tasks. Table 1 of Appendix C presents basic statistics for these datasets. For each dataset, we transform the explicit ratings into implicit feedback. To improve data quality, we employ filtering criteria, removing records with fewer than 10 interactions for all users and items, in accordance with existing research [40, 41].
IV-A2 Parameter Settings and Evaluation.
@@
We obtain optimal hyperparameters by optimizing the loss function (10) using the Adam optimizer with a learning rate of 0.001. The weight $\alpha$ for the prototype-based CL losses is set to 0.01. Additionally, the embedding size is fixed at 64, and the batch size is set to 128. The temperature coefficient in CL is established at 0.2, and the cluster number is set to 10. Furthermore, we apply batch normalization, dropout, and early stopping techniques to prevent overfitting. We use Hit Ratio (HR) and Normalized Discounted Cumulative Gain (NDCG) as evaluation metrics, which are frequently used in CDR methods [42, 3, 5].
IV-A3 Baseline Methods.
We compare the performance of FedPCL-CDR with SOTA baselines including NeuMF [43], LightGCN [34], FedNCF [44], GA-DTCDR [7], NMCDR [10], CL-DTCDR [11], GA-MTCDR-P [5], PriCDR [13], and P2FCDR [14], which are widely used in CDR approaches [24, 11, 9]. In Appendix D, we provide detailed descriptions of these baselines.
@@
TABLE I: Experimental results on four CDR tasks. The best performance is in bold, and the second best is underlined.
@@
Task Domain Metric NeuMF LightGCN FedNCF GA-DTCDR GA-MTCDR-P NMCDR CL-DTCDR PriCDR P2FCDR FedPCL-CDR Imp
Task 1 Phone HR 0.3490 0.3552 0.3477 0.3809 0.3827 0.3841 0.5277 0.3553 0.3756 0.5933 6.56%
NDCG 0.2141 0.2267 0.2034 0.2285 0.2304 0.2349 0.3193 0.2259 0.2267 0.3949 7.56%
Sport HR 0.3134 0.3257 0.3018 0.3749 0.3799 0.3914 0.5660 0.3687 0.4235 0.6766 11.06%
NDCG 0.1886 0.1939 0.1593 0.2056 0.2148 0.2166 0.3344 0.2389 0.2578 0.4534 11.90%
Task 2 Elec HR 0.4065 0.4124 0.3921 0.4568 0.4585 0.5234 0.5404 0.4234 0.4873 0.6078 6.74%
NDCG 0.3300 0.3315 0.3223 0.3402 0.3417 0.3316 0.3592 0.3534 0.3552 0.4018 4.26%
Cloth HR 0.2207 0.2367 0.2133 0.3044 0.2969 0.3030 0.4435 0.2816 0.3021 0.5018 5.83%
NDCG 0.1215 0.1402 0.1045 0.2156 0.2024 0.1965 0.2511 0.1856 0.2035 0.3166 6.55%
Phone HR 0.3490 0.3552 0.3477 0.4021 0.4093 0.4823 0.5159 0.3915 0.4329 0.5970 8.11%
NDCG 0.2141 0.2267 0.2034 0.2614 0.2705 0.3142 0.3059 0.2621 0.2804 0.3844 7.02%
Task 3 Movie HR 0.3294 0.3325 0.3145 0.3629 0.3691 0.3787 0.3525 0.3459 0.3703 0.4379 8.54%
NDCG 0.1768 0.1867 0.1505 0.1955 0.2014 0.2156 0.2016 0.2286 0.2037 0.2356 3.40%
Music HR 0.2686 0.2639 0.2987 0.3142 0.3187 0.3225 0.3351 0.3044 0.3223 0.4518 11.67%
NDCG 0.1468 0.1503 0.1789 0.1980 0.2004 0.2288 0.2310 0.2072 0.2159 0.2487 1.77%
Task 4 Book HR 0.2907 0.3045 0.2847 0.3869 0.3891 0.3796 0.4431 0.3525 0.3684 0.4724 2.93%
NDCG 0.1830 0.1904 0.1809 0.2260 0.2318 0.2316 0.2772 0.1902 0.2132 0.2865 0.93%
Movie HR 0.3294 0.3325 0.3145 0.374 0.3687 0.3703 0.4328 0.3336 0.3856 0.4601 2.73%
NDCG 0.1768 0.1867 0.1505 0.1894 0.1769 0.1968 0.2221 0.1530 0.2046 0.2444 2.23%
Music HR 0.2686 0.2639 0.2987 0.3142 0.3178 0.3525 0.3521 0.3050 0.3169 0.4745 12.20%
NDCG 0.1468 0.1503 0.1689 0.1780 0.1795 0.1888 0.1877 0.1685 0.1736 0.2059 1.71%
IV-B Experimental Results and Analysis
IV-B1 Performance Evaluation.

### v2

@@ 3.3.2 Differential Prototype Learning
These prototypes denote generalized user preferences, making it difficult for attackers to infer individual user details. To further improve privacy protection, we apply the LDP technique to the representative prototypes before transferring them. Specifically, we generate differential prototypes by adding Laplace noise to the representative prototypes:
$\hat{\textbf{c}}_{j}^{i}=clip(\textbf{c}_{j}^{i},\beta)+Lap(0,\eta);\hat{C}^{i}=\{\hat{\textbf{c}}_{j}^{i}\}_{j=1}^{K^{\prime}},$ (5)
where $\beta$ is the clipping threshold, and $\eta$ is the standard deviation of Laplace distribution. The LDP mechanism ensures that the leakage of private information is bounded by a privacy budget $\epsilon$, which has an upper bound $\frac{2\beta}{\eta}$, as derived in (Wu et al., 2022). A smaller privacy budget improves privacy protection but sacrifices the model’s performance. In our experimental evaluation, we empirically examine the utility-privacy trade-off of LDP through the comprehensive analysis of the privacy budget.
@@
Finally, we upload each domain’s differential prototype set $\hat{C}^{i}$ and overlapping user set $O^{i}$ to the global server.
3.3.3 Prototype-based Contrastive Learning
The server aggregates differential prototypes from all domains to generate both local and global prototypes (as detailed in the next section) and then distributes them back to the local domains.
To effectively transfer knowledge from other domains to the domain $D^{i}$, we enforce the alignment of user embeddings with corresponding global prototypes while distancing them from distinct global prototypes. The global prototype-based CL loss is defined as follows:
$\small L_{global}^{i}=-log\frac{exp(f_{s}(\textbf{e}_{u}^{i},\textbf{g}_{k}^{i}))}{exp(f_{s}(\textbf{e}_{u}^{i},\textbf{g}_{k}^{i}))+\sum_{\textbf{g}_{j}^{i}\in A(\textbf{g}_{k}^{i}),j\not=k}exp(f_{s}(\textbf{e}_{u}^{i},\textbf{g}_{j}^{i}))},$ (7)
where $\textbf{g}_{k}^{i}$ denotes the global prototype corresponding to cluster $k$, to which the user embedding $\textbf{e}_{u}^{i}$ belongs. We regard $(\textbf{e}_{u}^{i},\textbf{g}_{k}^{i})$ that belongs to the same cluster as a positive pair. Conversely, $\textbf{g}_{j}^{i}$ denotes the global prototype corresponding to cluster $j$, to which user embedding $\textbf{e}_{u}^{i}$ doesn’t belong, and $(\textbf{e}_{u}^{i},\textbf{g}_{j}^{i})$ forms a negative pair. $A(\textbf{g}_{k}^{i})$ is the set of global prototypes excluding $\textbf{g}_{k}^{i}$. $f_{s}$ indicates a similar function. We define it as:
$f_{s}(\textbf{e}_{u}^{i},\textbf{g}_{k}^{i})=\frac{\textbf{e}_{u}^{i}\cdot\textbf{g}_{k}^{i}}{||\textbf{e}_{u}^{i}||\times||\textbf{g}_{k}^{i}||}/\tau,$ (8)
where $\tau$ represents the temperature coefficient, which controls the concentration strength of representation (Wang and Liu, 2021).
Furthermore, we introduce the local prototype-based CL loss to align $\textbf{e}_{u}^{i}$ with local prototypes of each domain through domain-wise CL in the latent space, enhancing inter-domain knowledge sharing. It is defined as:
@@ 3.3.3 Prototype-based Contrastive Learning
The global and local prototypes capture cluster-relevant information at different granularities, guiding the transfer of user interests from various perspectives.
3.3.4 Local Training
After refining the user embedding $\textbf{e}_{u}^{i}$, we concatenate it with the item embedding $\textbf{e}_{v}^{i}$ and process them through MLP layers for preference prediction. The objective is to minimize the following loss:
@@ 3.3.4 Local Training
where $l$ is the cross-entropy loss function, $\hat{r}^{i}$ and $r^{i}$ are the predictive and ground-truth labels, respectively. The total loss is:
@@
where $\alpha$ is the trade-off parameter that balances $L_{prd}^{i}$ and prototype-based CL losses. The training procedure is detailed in Algorithm 1.
Algorithm 1 FedPCL-CDR
0:  $D^{i},\Theta_{i},i=1,2,...,M$
1:  Server executes:
2:  Initialize global prototype set $G=\{G^{1},G^{2},...,G^{M}\}$ and local prototype set $L=\{L^{1},L^{2},...,L^{M}\}$.
3:  for round $r=1$ to $R$ do
4:     for domain $i=1$ to $M$ do
5:        $O^{i},\hat{C}^{i}\leftarrow LocalUpdate(i,G^{i},L^{i})$
6:     end for
7:     Update global and local prototypes by Eq. (13) and Eq. (14).
8:  end for
9:  LocalUpdate($i,G^{i},L^{i},$)
10:  for epoch $e=1$ to $E$ do
11:     for batch $b=1$ to $B$ do
12:        Compute $L_{prd}^{i}$ by Eq. (10).
13:        if length($G^{i}$)==0 and length($L^{i}$)==0 then
14:           $L_{global}^{i}=0$, $L_{local}^{i}=0$
15:        else
16:           Compute $L_{global}^{i}$ by Eq. (7).
17:           Compute $L_{local}^{i}$ by Eq. (9).
18:        end if
19:        Update $\Theta_{i}$ by Eq. (11).
20:     end for
21:  end for
22:  Calculate prototype set $T^{i}$ by Eq. (3).
23:  Obtain representative prototype set $C^{i}$ by Eq. (4).
24:  Obtain differential prototype set $\hat{C}^{i}$ by Eq. (5).
25:  Obtain overlapping user set $O^{i}$ by Eq. (6).
26:  return  $O^{i}$ and $\hat{C}^{i}$
3.4 Global Server Aggregation
After receiving differential prototype set $\hat{C}$ and overlapping user set $O$ from all domains, the global server calculates global prototypes for each domain. First, for overlapping user set $o_{k}^{i}\in O^{i}$, we construct the prototype set $\bar{C}_{k}^{i}$ that includes all differential prototypes containing overlapping users with $o_{k}^{i}$ across domains:
$\bar{C}_{k}^{i}=\bigcup_{{i^{\prime}}\in D,{k^{\prime}}\leq{K^{\prime}}}\{\hat{\textbf{c}}_{k^{\prime}}^{i^{\prime}}|o_{k^{\prime}}^{i^{\prime}}\cap o_{k}^{i}\not=\emptyset\}.$ (12)
@@ 3.4 Global Server Aggregation
$\textbf{g}_{k}^{i}=\frac{1}{\hat{K}}\sum_{j=1}^{\hat{K}}\hat{\textbf{c}}_{j}^{i},$ (13)
where $\hat{K}$ is the length of prototype set $\bar{C}^{i}_{k}$.
@@
Different from the global prototypes, for the local prototypes, we select some differential prototypes in all domains via similarity calculation. Specifically, for each $\hat{\textbf{c}}_{k}^{i}$, we first calculate the cosine similarity between $\hat{\textbf{c}}_{k}^{i}$ and differential prototypes from other domains that include overlapping users with $o_{k}^{i}$. Then, we select the differential prototype with maximum similarity to $\textbf{c}_{k}^{i}$ in each domain to form the local prototype set $L_{k}^{i}$ for cluster $k$:
@@
3.5 Privacy Preserving Analysis
@@
The proposed FedPCL-CDR ensures user privacy through multiple mechanisms. Firstly, within the FL framework, data in each domain remains localized and is never shared with other domains, significantly reducing privacy risks (Wu et al., 2022). Secondly, cross-domain knowledge transfer is facilitated through prototypes, which inherently protect data privacy (Tan et al., 2022). These prototypes are one-dimensional vectors derived from averaging low-dimensional representations of samples within the same group, making the process irreversible. Lastly, we incorporate LDP on prototypes before transferring them, ensuring that the privacy leakage is bounded and further strengthening privacy protection (Qi et al., 2020).
3.6 Time Complexity Analysis
The total time complexity includes two main components: local domain learning and global server aggregation. For the local domain, the time complexity of graph learning is $O(|U^{i}|+|V^{i}|+|Y|)dL$, where $|Y|$ is the number of user-item interactions, $d$ is the embedding dimension, and $L$ is the number of layers. The time complexity of k-means is $O(|U^{i}|KTd)$ where T is the number of iterations. The time complexity of prototype-based CL is $O(|U^{i}|(|L^{i}|+|G^{i}|)d)$. The time complexity of MLP layers is $O(|Y|dH)$ where $H$ is the number of MLP layers. For the global server aggregation, the time complexity is $O(K^{\prime}{|\hat{C}_{k}^{i}|}^{2}+K^{\prime}|\hat{C}_{k}^{i}|)$. Therefore, the overall time complexity is $M((|U^{i}|+|V^{i}|+|Y|)dL+|U^{i}|KTd+|U^{i}|(|L^{i}|+|G^{i}|)d+|Y|dH)+K^{\prime}{|\hat{C}_{k}^{i}|}^{2}+K^{\prime}|\hat{C}_{k}^{i}|$. Due to $|U^{i}|+|V^{i}|<<|Y|$ and $|\hat{C}_{k}^{i}|<<|{\hat{C}_{k}^{i}|}^{2}$, The final time complexity can be simplified to $M(((L+H)|Y|+|U^{i}|(KT+|L^{i}|+|G^{i}|))d)+K^{\prime}{|\hat{C}_{k}^{i}|}^{2}$.
4 Experiments
We evaluate FedPCL-CDR through extensive experiments across four CDR tasks constructed from the Amazon and Douban datasets. Our goal is to address the following research questions:
  • 1.
    RQ1: How does FedPCL-CDR compare to state-of-the-art baselines in CDR tasks?
  • 2.
    RQ2: What is the contribution of each key component to FedPCL-CDR’s performance?
  • 3.
    RQ3: Does FedPCL-CDR address the sparse user overlap problem and ensure privacy protection?
  • 4.
    RQ4: How sensitive is FedPCL-CDR to critical hyperparameters?
4.1 Experimental Settings
4.1.1 Datasets
@@
Motivated by CDR methods (Chen et al., 2023a; Zhang et al., 2023; Cao et al., 2022; Liu et al., 2025), we conduct experiments on four domains from the Amazon dataset [^1] https://cseweb.ucsd.edu/ jmcauley/datasets/amazon/links.html: Cell Phones and Accessories (Phone), Electronics (Elec), Clothing, Shoes and Jewelry (Cloth), and Sports and Outdoors (Sport), and three domains from the Douban dataset [^2] https://www.dropbox.com/s/u2ejjezjk08lz1o/Douban.tar.gz?e=2&dl=0: Book, Movie, and Music. These are organized into four CDR tasks. Basic dataset statistics are summarized in Table 1. To align with implicit feedback settings, we binarize explicit ratings (0–5) into binary labels (1 or 0), consistent with the approach in (Acharyya and Pervin, 2025). Following prior studies (Kang et al., 2019; Zhao et al., 2020), we apply standard filtering to enhance data quality, discarding users and items with fewer than 10 interactions.
Table 1: Dataset statistics.
Tasks Datasets # Users # Items # Overlapping Users # Interactions Density
Task 1 Phone 5730 22287 655 82111 0.064%
Sport 10849 35368 172241 0.045%
Task 2 Elec 12301 56081 468 253300 0.037%
Cloth 13058 62137 185551 0.023%
Phone 5730 22287 82111 0.064%
Task 3 Movie 2320 5803 1134 102864 0.764%
Music 1193 7146 76592 0.898%
Task 4 Book 1715 8660 1008 104537 0.703%
Movie 2320 5803 102864 0.764%
Music 1193 7146 76592 0.898%
4.1.2 Evaluation Metrics
Motivated by previous CDR practice (Liu et al., 2020; Zhu et al., 2023b), we evaluate model performance via the leave-one-out approach. For each user, one interaction is randomly selected to form the test set, while the rest are used for training. Following (He et al., 2017; Wang et al., 2025), we sample 99 uninteracted items as negative samples for each test user, with the actual user-item interaction serving as the positive sample. The FedPCL-CDR model then generates prediction scores for all 100 candidates to facilitate ranking. Evaluation is conducted using Hit Ratio (HR) and Normalized Discounted Cumulative Gain (NDCG), which are frequently used in CDR methods (Li and Tuzhilin, 2020; Zhu et al., 2019, 2021; Ni et al., 2024).
4.1.3 Parameter Settings
@@
We obtain optimal hyperparameters by optimizing the loss function (11) using the Adam optimizer. The learning rate is set to 0.001. The weight $\alpha$ for the prototype-based CL losses is set to 0.01. Additionally, embedding vectors are 64-dimensional, and training is performed with a batch size of 256. The temperature coefficient in CL is established at 0.2, and the cluster number is set to 10. Furthermore, we apply batch normalization, dropout, and early stopping techniques to prevent overfitting. All baselines are executed using their GitHub source code, with hyperparameters carefully tuned on our datasets for the best performance. Experiments are conducted using two NVIDIA Testra V100 (32G) GPUs.
4.1.4 Baseline Methods.
We evaluate FedPCL-CDR against several SOTA baselines, which are commonly adopted in recent CDR research (Lu et al., 2023; Lin et al., 2024).
  • 1.
    NeuMF (He et al., 2017) fuses collaborative filtering with neural networks to model user-item interactions.
  • 2.
    LightGCN (He et al., 2020b) simplifies GCN by propagating user and item embeddings directly through the interaction graph, avoiding complex operations or auxiliary data.
  • 3.
    FedNCF (Perifanis and Efraimidis, 2022) introduces FL into the model Neural Collaborative Learning (NCF) to protect user’s privacy.
  • 4.
    GA-DTCDR (Zhu et al., 2020) proposes a graphical and attentional framework to learn comprehensive user and item embeddings.
  • 5.
    NMCDR (Xu et al., 2023) is a neural node matching framework that conducts intra-knowledge and inter-knowledge fusion to improve the model performance.
  • 6.
    CL-DTCDR (Lu et al., 2023) introduces both intra-domain and inter-domain CL tasks to address the data-sparsity issue and facilitate efficient knowledge transfer.
  • 7.
    GA-MTCDR-P (Zhu et al., 2021) is an enhanced model of GA-DTCDR, which incorporates graph neural networks and multi-domain attention mechanisms to address the problem of negative transfer in multi-target domain recommendations.
  • 8.
    PriCDR (Chen et al., 2022) is a PPCDR framework that leverages DP to release the source domain’s rating matrix, which is subsequently transferred to the target domain.
  • 9.
    P2FCDR (Chen et al., 2023a) is a federated CDR model that incorporates LDP to safeguard user embeddings during inter-domain knowledge transfer.
  • 10.
    ${\textbf{P}^{\textbf{2}}}$ (Chen et al., 2023b) introduces a dual-target CDR framework that utilizes a private set intersection algorithm and prototype-based FL to protect user privacy.
@@
Table 2: Experimental results on four CDR tasks. The best performance is in bold, and the second best is underlined. All improvements are significant over baselines (t-test with $p<0.01$).
@@
Domain Metric NeuMF LightGCN FedNCF GA- DTCDR GA- MTCDR-P NMCDR CL- DTCDR PriCDR P2FCDR $P^{2}DTR$ FedPCL -CDR Imp
Phone HR 0.3490 0.3552 0.3477 0.3809 0.3827 0.3841 0.5277 0.3553 0.3756 0.3843 0.5501 2.24%
NDCG 0.2141 0.2267 0.2034 0.2285 0.2304 0.2349 0.3193 0.2259 0.2267 0.2401 0.3541 3.48%
Sport HR 0.3134 0.3257 0.3018 0.3749 0.3799 0.3914 0.5660 0.3687 0.4235 0.4326 0.5991 3.31%
NDCG 0.1886 0.1939 0.1593 0.2056 0.2148 0.2166 0.3344 0.2389 0.2578 0.2645 0.3812 4.68%
Elec HR 0.4065 0.4124 0.3921 0.4568 0.4585 0.5234 0.5204 0.4234 0.4873 0.4987 0.5408 1.74%
NDCG 0.3000 0.3015 0.2923 0.3102 0.3117 0.3016 0.3292 0.3234 0.3252 0.3325 0.3445 1.20%
Cloth HR 0.2207 0.2367 0.2133 0.3044 0.2969 0.3030 0.4235 0.2816 0.3021 0.3124 0.4489 2.54%
NDCG 0.1215 0.1402 0.1045 0.2156 0.2024 0.1965 0.2511 0.1856 0.2035 0.2201 0.2678 1.67%
Phone HR 0.3490 0.3552 0.3477 0.4021 0.4093 0.4823 0.5159 0.3915 0.4329 0.4451 0.5341 1.82%
NDCG 0.2141 0.2267 0.2034 0.2614 0.2705 0.3142 0.3059 0.2621 0.2804 0.2943 0.3417 2.75%
Movie HR 0.3294 0.3325 0.3145 0.3629 0.3691 0.3787 0.3525 0.3459 0.3703 0.3825 0.4336 5.11%
NDCG 0.1768 0.1867 0.1505 0.1955 0.2014 0.2156 0.2016 0.2186 0.2037 0.2201 0.2301 1.00%
Music HR 0.2686 0.2639 0.2987 0.3142 0.3187 0.3025 0.3151 0.3044 0.3023 0.3156 0.3294 1.07%
NDCG 0.1468 0.1503 0.1589 0.1680 0.1704 0.1788 0.1631 0.1772 0.1659 0.1703 0.1842 0.54%
Book HR 0.2907 0.3045 0.1847 0.3069 0.3091 0.3096 0.3431 0.2525 0.2684 0.2764 0.3727 2.96%
NDCG 0.1830 0.1904 0.1409 0.1826 0.1918 0.2016 0.2072 0.1902 0.1932 0.2031 0.2265 1.93%
Movie HR 0.3294 0.3325 0.3145 0.3740 0.3687 0.3703 0.4328 0.3336 0.3856 0.3902 0.4470 1.42%
NDCG 0.1768 0.1867 0.1505 0.1894 0.1769 0.1968 0.2221 0.1530 0.2046 0.2135 0.2334 1.13%
Music HR 0.2686 0.2639 0.2987 0.3142 0.3178 0.3225 0.3121 0.3050 0.3069 0.3124 0.3418 1.93%
NDCG 0.1468 0.1503 0.1689 0.1780 0.1795 0.1888 0.1877 0.1685 0.1736 0.1735 0.2034 0.47%
4.2 Experimental Results and Analysis
4.2.1 Performance Evaluation (RQ1)